A 0.38-kg skeet (clay target) is fired at an angle of 28∘ to the horizon with a speed of 25 m/s. When it reaches the maximum height, h, it is hit from below by a 28-g pellet traveling vertically upward at a speed of 230 m/s. The pellet is embedded in the skeet.
How much higher, h′, did the skeet go up?
Express your answer to two significant figures and include the appropriate units.
1
Expert's answer
2019-04-01T10:40:35-0400
"h=\frac{v^2\sin^2{28}}{2g}=\frac{25^2\sin^2{28}}{2(9.8)}=7 m."

From the conservation of momentum


"mu=(m+M)v'"

From the conservation of energy:


"0.5(m+M)v'^2=(m+M)gH"

"H=\frac{v'^2}{2g}=\frac{u^2m^2}{2(m+M)^2g}=\frac{230^2(0.028)^2}{2(0.028+0.38)^2 (9.8)}=13 m."

So,


"h'=h+H=7+13=20 m."